### v2

@@ Abstract
Inertia of a particle is due to its mass as well as intrinsic spin. The latter is revealed via the coupling of intrinsic spin with rotation. The spin-rotation coupling and the concomitant spin-gravity coupling are discussed in connection with the nature of inertia. The spin-rotation-gravity coupling leads to a gravitomagnetic Stern-Gerlach type of force on the particle that is independent of the particle’s mass and thus violates the universality of free fall. This effect is extremely small and its measurement is far beyond present capabilities.
@@ II.1 Spin-Rotation Coupling
To extend the theory of relativity beyond the WKB limit has been the aim of the nonlocal special relativity theory Mashhoon:1997qc ; BM5 ; Mashhoon:2008vr . A comprehensive treatment of spin-rotation coupling and a more complete list of references are contained in BMB , while the results of recent spin-rotation experiments in connection with neutron physics can be found in DSH ; DDSH ; DDKWLSH .
@@ III Spin-gravity coupling
The intrinsic spin-gravity coupling has been the subject of numerous investigations; see DeOT ; BM ; Bah1 ; Bah2 ; Ramos:2006sb ; Mashhoon:2008si ; Papini:2007gx ; Mashhoon:2013jaa ; Tarallo:2014oaa ; Fadeev:2020gjk ; Mashhoon:2023idh ; Vergeles:2022mqu and the references cited therein.
@@ IV Gravitomagnetic Stern-Gerlach Effect
which are independent of the inertial mass and thus violate the universality of free fall. In the correspondence limit, the gravitomagnetic Stern-Gerlach force is in agreement with the classical Mathisson spin-curvature force Mashhoon:2021qtc .
@@ VI DISCUSSION
We have argued that there is a direct coupling between the intrinsic spin of a particle and rotation as a consequence of the inertia of intrinsic spin. Accelerated motion consists in general of translational acceleration as well as rotation. Is there a direct coupling between intrinsic spin and translational acceleration as well? A direct interaction Hamiltonian of the form $\bm{S}\cdot\bm{a}/c$ or $\bm{S}\cdot\bm{g}/c$ would violate parity and time-reversal invariance; here, $\bm{a}$ represents the translational acceleration and $\bm{g}$ represents the acceleration of gravity. The $\bm{S}\cdot\bm{g}/c$ term has been ruled out observationally Bah2 ; W+R . Furthermore, no such direct spin-acceleration coupling has been found in the study of wave equations in translationally accelerated systems in Minkowski spacetime and in the corresponding gravitational fields Mashhoon:2004rz ; Bini:2004kz .

### v3

@@ Abstract
Inertia of a particle is due to its mass as well as intrinsic spin. The latter is revealed via the coupling of intrinsic spin with rotation. The spin-rotation coupling and the concomitant spin-gravity coupling are discussed in connection with the nature of inertia. The spin-rotation-gravity coupling leads to a gravitomagnetic Stern-Gerlach type of force on the particle that is independent of the particle’s mass and thus violates the universality of free fall. This effect is extremely small and its measurement is beyond present capabilities.
@@ II.1 Spin-Rotation Coupling
To extend the theory of relativity beyond the WKB limit has been the aim of the nonlocal special relativity theory Mashhoon:1997qc ; BM5 ; Mashhoon:2008vr . A comprehensive treatment of spin-rotation coupling and a more complete list of references are contained in BMB , while the results of recent spin-rotation experiments in connection with neutron physics can be found in DSH ; DDSH ; DDKWLSH ; Geerits:2024jdt .
@@ III Spin-gravity coupling
The intrinsic spin-gravity coupling has been the subject of numerous investigations; see DeOT ; BM0 ; BM ; Bah1 ; Bah2 ; Ramos:2006sb ; Mashhoon:2008si ; Papini:2007gx ; Mashhoon:2013jaa ; Tarallo:2014oaa ; Fadeev:2020gjk ; Mashhoon:2023idh ; Vergeles:2022mqu and the references cited therein. A comprehensive account of spin-gravity interaction is contained in Lambiase:2021txu .
@@ IV Gravitomagnetic Stern-Gerlach Effect
which are independent of the particle’s mass and thus violate the universality of free fall. In the correspondence limit, the gravitomagnetic Stern-Gerlach force is in agreement with the classical Mathisson spin-curvature force Mashhoon:2021qtc .
@@ VI DISCUSSION
We have argued that there is a direct coupling between the intrinsic spin of a particle and rotation as a consequence of the inertia of intrinsic spin. Accelerated motion consists in general of translational acceleration as well as rotation. Is there a direct coupling between intrinsic spin and translational acceleration as well? A direct interaction Hamiltonian of the form $\bm{S}\cdot\bm{a}/c$ or $\bm{S}\cdot\bm{g}/c$ would violate parity and time-reversal invariance; here, $\bm{a}$ represents the translational acceleration and $\bm{g}$ represents the acceleration of gravity. The $\bm{S}\cdot\bm{g}/c$ term has been ruled out observationally Bah2 ; W+R . Furthermore, no such direct spin-acceleration coupling has been found in the study of wave equations in translationally accelerated systems in Minkowski spacetime and in the corresponding gravitational fields Bah3 ; Mashhoon:2004rz ; Bini:2004kz .